OpenStudy (anonymous):

You are given the following function. f(t) = t + √t Find the derivative of the function using the definition of derivative.

OpenStudy (anonymous):

\[\begin{align*}f'(t)&=\lim_{h\to0}\frac{f(t+h)-f(t)}{h}\\ &=\lim_{h\to0}\frac{(t+h+\sqrt{t+h})-(t+\sqrt t)}{h}\\ &=\lim_{h\to0}\frac{h+\sqrt{t+h}-\sqrt t}{h}\\ &=\lim_{h\to0}\frac{h}{h}+\lim_{h\to0}\frac{\sqrt{t+h}-\sqrt t}{h}\\ &=\lim_{h\to0}1+\lim_{h\to0}\frac{\sqrt{t+h}-\sqrt t}{h}\\ &=1+\lim_{h\to0}\frac{\sqrt{t+h}-\sqrt t}{h} \end{align*}\] What do you think you can do about those square roots?

OpenStudy (anonymous):

Try multiplying by the conjugate